OpenStudy (anonymous):

Evaluate f(-2) if f(x)=(1+x/1-x)^2

OpenStudy (eyust707):

f(-2) means plug -2 in for x

OpenStudy (anonymous):

put x=-2 into f(x)=(1+x/1-x)^2. You can get the answer f(-2) = ( 1-2 /1-(-2))^2

jhonyy9 (jhonyy9):

f(-2)=1/9

OpenStudy (anonymous):

would it be -1/9?

jhonyy9 (jhonyy9):

no

jhonyy9 (jhonyy9):

because -1 squared is plus 1

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

so the answer would be 1 correct?

jhonyy9 (jhonyy9):

yes 1/9

OpenStudy (anonymous):

i dont have that as one of the chocies though i have -1/9, 1, -1, and 3

jhonyy9 (jhonyy9):

so -1/9 can being just if when the numerator not will be squared

jhonyy9 (jhonyy9):

but because you have wrote in your exercise all fraction squared so hence not can being the result minus

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

check it your exercise how is write in your book